# Negate or Embrace: On How Misalignment Shapes Multimodal Representation Learning

### v3

@@ B.1 Lemmas
    Finally, we demonstrate that the optimized mappings are invertible with respect to the learned representations and the true unbiased semantic variables $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$. This ensures that the representations block-identify the preserved unbiased semantic variables, thereby concluding the proof.
Step 1
(Global Minimum and Invariance Condition). Let $\mathbf{f}_{x}:\mathcal{X}\to(0,1)^{n}$ and $\mathbf{f}_{t}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{n}$ be any smooth functions attaining the global minimum. Define the smooth mappings:
@@
Since all terms in $\mathcal{L}_{\text{SymAlignMaxEnt}}$ are non-negative, and its global minimum is 00 by Lem. B.1, each term in $\mathcal{L}_{\text{SymAlignMaxEnt}}$ must vanish a.s. for any pairing $(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})$, leading to:
@@
$\mathbf{h}_{t}(\mathbf{z}_{t^{\scriptscriptstyle(\theta)}})$ $=\mathbf{h}_{x}(\mathbf{z}_{x}),\quad\text{a.s.}\quad\forall\,\mathbf{z}_{x}\sim p_{\mathbf{s}}\,p_{\mathbf{m}_{x}},\;\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}\sim p_{\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\mid\mathbf{s},\theta,\rho}\,p_{\mathbf{m}_{t}},$ (12)
@@
Since $\mathbf{z}_{x}=(\mathbf{s},\mathbf{m}_{x})$ and $\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}=(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t})$ are the latent variables generating images $\mathbf{x}$ and paired text $\mathbf{t}^{\scriptscriptstyle(\theta)}$, respectively, the assumed data-generating process (§ 3) implies the following independence properties:
@@
A symmetric argument applies to $\mathbf{m}_{t}$. If $\mathbf{m}_{t}\mapsto\mathbf{m}_{t}+\boldsymbol{\varsigma}$, then by Eq. 14 in statement (c2), $\mathbf{h}_{x}(\mathbf{s},\mathbf{m}_{x},\mathbf{m}_{t}+\boldsymbol{\varsigma})$ remains unchanged. The invariance condition in Eq. 12 then forces $\mathbf{h}_{t}(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t}+\boldsymbol{\varsigma})$ to remain constant w.r.t. $\boldsymbol{\varsigma}$, showing that $\mathbf{h}_{t}$ is independent of $\mathbf{m}_{t}$. Therefore, the learned representations satisfy:
@@
Using the invariant condition in Eq. 12, together with the independence of modality-specific non-semantic variables in Equations 17 and 18, we have the following updated invariant condition:
@@
$\exists\,l\in\mathbb{I}^{c}_{\theta},\quad(\mathbf{s}^{*}_{\mathbb{I}_{\theta}},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\theta}})\in\mathcal{S},\quad\text{such that}\quad\frac{\partial\,\mathbf{h}^{c}_{x}\bigl{(}\mathbf{s}^{*}_{\mathbb{I}_{\theta}},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\theta}}\bigr{)}}{\partial\,s^{*}_{l}}\;\neq\;0.$
@@
$\exists\,\eta>0:\quad s_{l}\;\mapsto\;\mathbf{h}^{c}_{x}\!\bigl{(}\mathbf{s}^{*}_{\mathbb{I}_{\theta}},(s_{l},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\theta}\setminus\{l\}})\bigr{)}\quad\text{is strictly monotonic on}\quad(s^{*}_{l}-\eta,\;s^{*}_{l}+\eta),$
@@
where $\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\theta}}$ represents the perturbed semantic variables of $\mathbf{s}^{*}_{\mathbb{I}_{\theta}}$ introduced by selection bias (with the exclusion of perturbed components further addressed below).
@@
(i) Justification for the existence of distinct points in Eq. 20. This follows from the assumption $p_{\mathbf{s}}>0$ a.e. on $\mathcal{S}$ by Asm. 3.1. From a latent SCM perspective (Pearl, 2009; von Kügelgen et al., 2024), even if a specific semantic component $s_{l}$ in $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ is the ancestor node of some other semantic components in $\mathbf{s}_{\mathbb{I}_{\theta}}$, the strict positivity of $p_{\mathbf{s}}$ ensures that the exogenous noise variables are well-defined. Thus, for different values of $s_{l}$, there exist corresponding noise values that keep $\mathbf{s}^{*}_{\mathbb{I}_{\theta}}$ remaining fixed. (ii) What if the unknown causal structure is $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}\to\mathbf{s}_{\mathbb{I}_{\theta}}$? The potential causal influence from $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ to $\mathbf{s}_{\mathbb{I}_{\theta}}$ does not resolve the contradiction. Independence here means that, once $\mathbf{s}_{\mathbb{I}_{\theta}}$ is set, there is no direct functional path from $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ to the representations $\mathbf{h}_{x}(\mathbf{s}_{\mathbb{I}_{\theta}})$, i.e., the causal influence among them is fully accounted for by the realized value of $\mathbf{s}_{\mathbb{I}_{\theta}}$.
@@
(Exclusion of Perturbed Semantic Variables). We now demonstrate that both representations are independent of $\mathbf{s}_{\mathbb{I}_{\rho}}$ and $\tilde{\mathbf{s}}_{\mathbb{I}_{\rho}}$ respectively, as a consequence of the contradiction between the invariant condition and the random perturbations introduced by perturbation bias.
First, we refine the invariant condition by excluding omitted semantic variables as established above. Combining Equations 12, 18 and 23, we obtain:
@@
$\exists\,l\in\mathbb{I}_{\rho},\quad(\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\rho}},\tilde{\mathbf{s}}^{*}_{\mathbb{I}^{c}_{\rho}}),\quad\text{such that}\quad\frac{\partial\mathbf{h}^{c}_{t}(\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\rho}},\tilde{\mathbf{s}}^{*}_{\mathbb{I}^{c}_{\rho}})}{\partial\tilde{s}^{*}_{l}}\neq 0.$
@@
On the other hand, by the pairing conditions in Asm. 3.2, there exists at least one subset $A\subseteq\mathbb{I}_{\rho}$ of perturbed semantic variables such that $l\in A$ and $p_{A}(A)>0$. Pick one such set and call it $A$. Define the latent semantic variables corresponding to the image of this pair as $(\mathbf{s}^{*}_{A},\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A})$ (here, we omit the components that have already been excluded in previous steps).
@@
Thus, with a positive probability guaranteed by the above conditional, for the image semantic variables $(\mathbf{s}^{*}_{A},\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A})$, we can construct two distinct realizations of perturbed semantic variables for generating different text:
$\bigl{(}(s_{l}^{\scriptscriptstyle(1)},\tilde{\mathbf{s}}^{*}_{A\setminus\{l\}}),\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A}\bigr{)},\;\;\bigl{(}(s_{l}^{\scriptscriptstyle(2)},\tilde{\mathbf{s}}^{*}_{A\setminus\{l\}}),\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A}\bigr{)},$
@@
$s_{l}^{\scriptscriptstyle(1)}\in(\tilde{s}^{*}_{l}-\eta_{2},\tilde{s}^{*}_{l}),\quad s_{l}^{\scriptscriptstyle(2)}\in(\tilde{s}^{*}_{l},\tilde{s}^{*}_{l}+\eta_{2})\quad\text{with}\quad\eta_{2}=\min(\eta,\eta_{1}).$
Based on the invariant condition established in Eq. 24, we have the following equalities:
$\mathbf{h}^{c}_{t}\bigl{(}(s_{l}^{\scriptscriptstyle(1)},\tilde{\mathbf{s}}^{*}_{A\setminus\{l\}}),\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A}\bigr{)}=\mathbf{h}^{c}_{t}\bigl{(}(s_{l}^{\scriptscriptstyle(2)},\tilde{\mathbf{s}}^{*}_{A\setminus\{l\}}),\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A}\bigr{)}=\mathbf{h}_{x}(\mathbf{s}^{*}_{A},\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A}).$
@@
Updating the invariant condition, and combining Equations 24 and 26, we obtain:
@@
By directly applying Lem. B.2, it follows that the learned representations $\hat{\mathbf{z}}_{x}$ (and also $\hat{\mathbf{z}}_{t}$) include all and only the information of the unaltered semantic components $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ almost surely, and that $\mathbf{h}_{x}$ (and similarly $\mathbf{h}_{t}$) is invertible. Consequently, the true modality-shared semantic variables $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ are block-identified by $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$.
@@ C.1 Detailed Experimental Setup
For independent semantic variables, $\Sigma_{\mathbf{s}}=I_{10}$ yields a factorized standard Gaussian, aligning with nonlinear ICA assumptions (Khemakhem et al., 2020; Sorrenson et al., 2020). For dependent semantics, we sample $\Sigma_{\mathbf{s}}$ from Wishart distribution $\mathcal{W}_{10}(I,10)$ (identity scale, 10 degrees of freedom), introducing latent dependencies as in (Von Kügelgen et al., 2021; Daunhawer et al., 2022; Yao et al., 2023). Modality-specific covariances $\Sigma_{\mathbf{m}_{x}}$ and $\Sigma_{\mathbf{m}_{t}}$ are sampled similarly from $\mathcal{W}_{5}(I,5)$, the Wishart distribution with identity scale and 5 degrees of freedom.
@@
In addition to isolated biases, we consider a joint bias scenario to investigate compounding effects. Specifically, we select $\mathbb{I}_{\theta}=[8]$ (corresponding $\theta=968$, excluding the last two semantic indices), and apply perturbations to the first two indices ($\mathbb{I}_{\rho}=\{1,2\}$, correspnding to $\rho=12$).
Results across these 20 representative configurations (10 selection, 10 perturbation) and the joint bias case serve to validate the theoretical findings under a range of misalignment settings.
@@
We parameterize the generative functions $\mathbf{g}_{x}$ and $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ using randomly initialized 3-layer invertible MLPs, following prior work (Daunhawer et al., 2022). Invertibility is enforced by maintaining a condition number threshold of $31\mathrm{e}{-}$ 3 for each layer.
@@
Each function $f_{y_{(\cdot)}}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ includes quadratic, cubic, pairwise, and triple-wise interaction terms, as well as sinusoidal, logarithmic, and exponential transformations. The full formulation for a semantic vector $\mathbf{s}_{[1:d]}$ is:

### v4

@@ B.1 Lemmas
    Finally, we demonstrate that the optimized mappings, which compose with the underlying generative functions, are invertible with respect to the learned representations and the true unbiased semantic variables $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$. This ensures that the representations block-identify the preserved unbiased semantic variables, thereby concluding the proof.
Step 1 (Global Minimum and Invariance Condition). Let $\mathbf{f}_{x}:\mathcal{X}\to(0,1)^{n}$ and $\mathbf{f}_{t}:\mathcal{T}^{\scriptscriptstyle(\theta)}\to(0,1)^{n}$ be any smooth functions attaining the global minimum. Define the smooth mappings:
@@
Since all terms in $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ are non-negative, and its global minimum is 00 by Lem. B.1, each term in $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ must vanish a.s. for any pairing $(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})$. Thus, minimizing $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ leads to:
@@
$\mathbf{h}_{t}(\mathbf{z}_{t^{\scriptscriptstyle(\theta)}})$ $=\mathbf{h}_{x}(\mathbf{z}_{x})\quad\text{a.s.}\quad\forall\,\mathbf{z}_{x}\sim p_{\mathbf{s}}\,p_{\mathbf{m}_{x}},\;\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}\sim p_{\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\mid\mathbf{s},\theta,\rho}\,p_{\mathbf{m}_{t}},$ (12)
@@
Since $\mathbf{z}_{x}=(\mathbf{s},\mathbf{m}_{x})$ and $\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}=(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t})$ are the latent variables generating images $\mathbf{x}$ and paired text $\mathbf{t}^{\scriptscriptstyle(\theta)}$, respectively, the data-generating process in § 3 implies the following independence properties:
@@
A symmetric argument applies to $\mathbf{m}_{t}$. If $\mathbf{m}_{t}\mapsto\mathbf{m}_{t}+\boldsymbol{\varsigma}$ with $\boldsymbol{\varsigma}$ a small enough perturbation, then by Eq. 14 in statement (c2), $\mathbf{h}_{x}(\mathbf{s},\mathbf{m}_{x},\mathbf{m}_{t}+\boldsymbol{\varsigma})$ remains unchanged. The invariance condition in Eq. 12 then forces $\mathbf{h}_{t}(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t}+\boldsymbol{\varsigma})$ to remain constant w.r.t. $\boldsymbol{\varsigma}$, showing that $\mathbf{h}_{t}$ is independent of $\mathbf{m}_{t}$. Therefore, the learned representations satisfy:
@@
Using the invariance condition in Eq. 12, together with the independence of modality-specific non-semantic variables in Equations 17 and 18, we have the following updated invariance condition:
@@
$\exists\,l\in\mathbb{I}^{c}_{\theta},\;(\mathbf{s}^{*}_{\mathbb{I}_{\theta}},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\theta}})\in\mathcal{S},\quad\text{such that}\quad\frac{\partial\,\mathbf{h}^{c}_{x}\bigl{(}\mathbf{s}^{*}_{\mathbb{I}_{\theta}},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\theta}}\bigr{)}}{\partial\,s^{*}_{l}}\;\neq\;0.$
@@
$\exists\,\eta>0:s_{l}\;\mapsto\;\mathbf{h}^{c}_{x}\!\bigl{(}\mathbf{s}^{*}_{\mathbb{I}_{\theta}},(s_{l},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\theta}\setminus\{l\}})\bigr{)}\quad\text{is strictly monotonic on}\quad(s^{*}_{l}-\eta,\;s^{*}_{l}+\eta),$
@@
where $\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\theta}}$ represents the perturbed semantic variables of $\mathbf{s}^{*}_{\mathbb{I}_{\theta}}$ introduced by selection bias (with the exclusion of perturbed components further addressed in the next step).
@@
(i) Justification for the existence of distinct points in Eq. 20. This follows from the assumption $p_{\mathbf{s}}>0$ a.e. on $\mathcal{S}$ by Asm. 3.1. From a latent SCM perspective [51, 65], even if a specific semantic component $s_{l}$ in $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ is the ancestor node of some other semantic components in $\mathbf{s}_{\mathbb{I}_{\theta}}$, the strict positivity of $p_{\mathbf{s}}$ ensures that the exogenous noise variables are well-defined. Thus, for different values of $s_{l}$, there exist corresponding noise values that keep $\mathbf{s}^{*}_{\mathbb{I}_{\theta}}$ remaining fixed. (ii) What if the unknown causal structure is $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}\to\mathbf{s}_{\mathbb{I}_{\theta}}$? The potential causal influence from $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ to $\mathbf{s}_{\mathbb{I}_{\theta}}$ does not resolve the contradiction. Independence here means that, once $\mathbf{s}_{\mathbb{I}_{\theta}}$ is set, there is no direct functional path from $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ to the representations $\mathbf{h}_{x}(\mathbf{s}_{\mathbb{I}_{\theta}})$, i.e., the causal influence among them is fully accounted for by the realized value of $\mathbf{s}_{\mathbb{I}_{\theta}}$.
@@
(Exclusion of Perturbed Semantic Variables). We now demonstrate that both representations are independent of $\mathbf{s}_{\mathbb{I}_{\rho}}$ and $\tilde{\mathbf{s}}_{\mathbb{I}_{\rho}}$ respectively, as a consequence of the contradiction between the invariance condition and the random perturbations introduced by perturbation bias.
First, we refine the invariance condition by excluding omitted semantic variables as established above. Combining Equations 12, 18 and 23, we obtain:
@@
$\exists\,l\in\mathbb{I}_{\rho},\;(\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\rho}},\tilde{\mathbf{s}}^{*}_{\mathbb{I}^{c}_{\rho}}),\quad\text{such that}\quad\frac{\partial\mathbf{h}^{c}_{t}(\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\rho}},\tilde{\mathbf{s}}^{*}_{\mathbb{I}^{c}_{\rho}})}{\partial\tilde{s}^{*}_{l}}\neq 0.$
@@
On the other hand, by the pairing conditions in Asm. 3.2, there exists at least one subset $A\subseteq\mathbb{I}_{\rho}$ of perturbed semantic variables such that $l\in A$ and $p_{A}(A)>0$. Pick one such set and call it $A$. Define the latent semantic variables corresponding to the image of this pair as $(\mathbf{s}^{*}_{A},\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A})$ (here, we omit the latent semantic components that have already been excluded in previous steps).
@@
Thus, with a positive probability guaranteed by the above conditional, for the image semantic variables $(\mathbf{s}^{*}_{A},\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A})$, we can construct two distinct realizations of perturbed semantic variables for generating different text (due to $p_{\mathbf{s}}>0$ a.e.):
$\bigl{(}(s_{l}^{\prime},\tilde{\mathbf{s}}^{*}_{A\setminus\{l\}}),\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A}\bigr{)},\;\;\bigl{(}(s_{l}^{\prime\prime},\tilde{\mathbf{s}}^{*}_{A\setminus\{l\}}),\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A}\bigr{)},$
@@
$s_{l}^{\prime}\in(\tilde{s}^{*}_{l}-\eta_{2},\tilde{s}^{*}_{l}),\quad s_{l}^{\prime\prime}\in(\tilde{s}^{*}_{l},\tilde{s}^{*}_{l}+\eta_{2})\quad\text{with}\quad\eta_{2}=\min(\eta,\eta_{1}).$
Based on the invariance condition established in Eq. 24, we have the following equalities:
$\mathbf{h}^{c}_{t}\bigl{(}(s_{l}^{\prime},\tilde{\mathbf{s}}^{*}_{A\setminus\{l\}}),\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A}\bigr{)}=\mathbf{h}^{c}_{t}\bigl{(}(s_{l}^{\prime\prime},\tilde{\mathbf{s}}^{*}_{A\setminus\{l\}}),\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A}\bigr{)}=\mathbf{h}_{x}(\mathbf{s}^{*}_{A},\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A}).$
@@
Updating the invariance condition, and combining Equations 24 and 26, we obtain:
@@
According to Equations 10 and 11, we know that the learned representations in both modalities are uniformly distributed over $(0,1)^{n}$. By directly applying Lem. B.2, it follows that the learned representations $\hat{\mathbf{z}}_{x}$ (and also $\hat{\mathbf{z}}_{t}$) include all and only the information of the unaltered semantic components $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ a.s., and that $\mathbf{h}_{x}$ (and similarly $\mathbf{h}_{t}$) is invertible. Consequently, the true modality-shared semantic variables $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ are block-identified by $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$.
@@ C.1 Detailed Experimental Setup
For independent semantic variables, $\Sigma_{\mathbf{s}}=I_{10}$ yields a factorized standard Gaussian, aligning with nonlinear ICA assumptions [26, 61]. For dependent semantics, we sample $\Sigma_{\mathbf{s}}$ from Wishart distribution $\mathcal{W}_{10}(I,10)$ (identity scale, 10 degrees of freedom), introducing latent dependencies as in [66, 12, 76]. Modality-specific covariances $\Sigma_{\mathbf{m}_{x}}$ and $\Sigma_{\mathbf{m}_{t}}$ are sampled similarly from $\mathcal{W}_{5}(I,5)$, the Wishart distribution with identity scale and 5 degrees of freedom.
@@
In addition to the isolated bias settings, we consider a joint bias scenario to investigate compounding effects. Specifically, we select $\mathbb{I}_{\theta}=[8]$ (corresponding $\theta=968$, excluding the last two semantic indices), and apply perturbations to the first two indices ($\mathbb{I}_{\rho}=[2]$, correspnding to $\rho=12$).
Results across these 20 representative configurations (10 selection settings, 10 perturbation settings) and the joint bias case serve to validate the theoretical findings under a range of misalignment settings.
@@
We parameterize the generative functions $\mathbf{g}_{x}$ and $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ using randomly initialized 3-layer invertible MLPs, following prior work [12]. Invertibility is enforced by maintaining a condition number threshold of $31\mathrm{e}{-}$ 3 for each layer.
@@
Each function $f_{y_{(\cdot)}}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ includes quadratic, cubic, pairwise, and triple-wise interaction terms, as well as sinusoidal, logarithmic, and exponential transformations. The full formulation for a semantic vector $\mathbf{s}_{[d]}$ is: